OpenStudy (anonymous):

how do you find all the solutions (real and imaginary) for this, 0=5x^3(cubed)-625

OpenStudy (anonymous):

I would first notice that you can factor out a 5 from both of those.

OpenStudy (anonymous):

So then you would have 0=5(x^3-125) 0=x^3-125, since you can divide 5 on both sides and 0/5 =0.

OpenStudy (anonymous):

Does this look similar to other types of problems you have for the functions to the third power?